PLEASE HELP
If the graph of f contains the points ( -1, 3), ( 0, 4) and ( 2, -7) list the points that would be found on the graphs of each of the following functions:
a) f (x – 3)
b) f (x) + 3
Thank You

Found 2 solutions by stanbon, shahid:
Answer by stanbon(75887)

If the graph of f contains the points ( -1, 3), ( 0, 4) and ( 2, -7) list the points that would be found on the graphs of each of the following functions:
a) f (x – 3)
Points move 3 units to the right:
(-1,3) goes to (2,3)
(0,4) goes to (3,4)
(2,-7) goes to (5,-7)
-----------------------------
b) f(x) + 3
Points move up 3
(-1,3) goes to (-1,6)
(0,4) goes to (0,7)
(2,-7) goes to (2,-4)
------------------------
Cheers,
Stan H.

Answer by shahid(44)

a)f(x-3)
from given points we get f(-1)=3 , f(0)=4 and f(2)=7
take x-3=-1 then x=2 so point (-1,3) corresponds to (2,3)
similarly (0,4) gives (3,4) and (2,7) gives (5,7) in this way we got these three points lying on the graph
b)f(x)+3 now consider h(x) such that
h(x)=f(x)+3 if x=-1 then
h(-1)=f(-1)+3
h(-1)= 3+3 since f(-1)=3
h(-1)= 6 so point is (-1,6) other points can be obtained by same